### v1

# Density-valued time series: Nonparametric density-on-density regression

Abstract
This paper is concerned with forecasting probability density functions. Density functions are nonnegative and have a constrained integral; they thus do not constitute a vector space. Implementing unconstrained functional time-series forecasting methods is problematic for such nonlinear and constrained data. A novel forecasting method is developed based on a nonparametric function-on-function regression, where both the response and the predictor are probability density functions. Through a series of Monte-Carlo simulation studies, we evaluate the finite-sample performance of our nonparametric regression estimator. Using French departmental COVID19 data and age-specific period life tables in the United States, we assess and compare finite-sample forecast accuracy between the proposed and several existing methods.
Keywords: Bayes space; Convex constraint; Density forecasting; Nonparametric function-on-function regression; Time series of densities.
Short Run Title: Nonparametric density-on-density regression
1 Introduction
Recent advances in computer storage and data collection have enabled researchers to record data of characteristics varying over a continuum (time, space, depth, wavelength, etc.). In diverse branches of science, the data are collected by a spectrometer, rain gauges, electroencephalographs, or even a high-performance computer. In all these cases, a number of subjects are observed densely over time, space, or both. Through the application of interpolation or smoothing techniques, these data become functions that can be represented as a curve, image, or shape. Functional data analysis has arisen as a field of statistics that provides statistical tools for analyzing this type of information. For an overview of functional data analysis, Ramsay & Silverman (2005) presented several state-of-the-art statistical techniques while Ferraty & Vieu (2006) listed a range of nonparametric techniques. Hsing & Eubank (2015) provided theoretical foundations with an introduction to linear operators, while Bosq (2000) studied functional autoregressive process from a linear operator perspective. An introduction to temporally dependent functional data can be found in Kokoszka & Reimherr (2017), while Mateu & Giraldo (2022) studied spatially-dependent functional data.
As an integral part of functional data analysis and time series analysis, functional time series consist of random functions observed at a time interval. Functional time series can be classified into two categories depending on whether the continuum is also a time variable. On the one hand, when the continuum is not a time variable, functional time series can also arise when observations in a period can be considered as finite-dimensional realisations of an underlying continuous function. Examples include yearly age-specific mortality rates (see, e.g., Chiou & Müller, 2009; Hyndman & Shang, 2009) and a time series of near-infrared spectroscopy curves (see, e.g., Shang et al., 2022). On the other hand, functional time series can arise from measurements obtained by separating a continuous-time stochastic process into natural consecutive intervals, for example, days, weeks, or years (see, e.g., Bosq, 2000; Besse et al., 2000; Antoniadis & Sapatinas, 2003; Antoniadis et al., 2006; Ferraty & Vieu, 2006; Hörmann & Kokoszka, 2012; Kokoszka et al., 2017; Shang, 2017).
Functional data that are samples of one-dimensional random probability density functions (PDFs) are common. Examples include income distributions (Kneip & Utikal, 2001), distributions of the times when bids are submitted in an online auction (Jank et al., 2008), functional connectivity in the brain (Petersen & Müller, 2019), distribution of image features from head CT scans (Salazar et al., 2020), distributions of stock return (Harvey et al., 2016; Bekierman & Gribisch, 2021; Zhang et al., 2022), the age distribution of fertility rates (Mazzuco & Scarpa, 2015) and age distribution of life-table death counts in demography (Bergeron-Boucher et al., 2017).
To model density-valued functional data, Jones & Rice (1992) estimated density function by kernel density estimator and displayed density function via principal components. Nerini & Ghattas (2007) proposed regression trees where the response variable is PDFs. van der Linde (2008) introduced Bayesian functional principal component analysis and applied it to simulated data consisting of nonparametric density estimates. Srivastava et al. (2011) proposed a time-warping function where the square root transformation of densities resides in the Hilbert space. Dai (2022) considered statistical inference of the Fréchet mean on the Hilbert sphere with application to random densities. Zhang et al. (2022) proposed Wasserstein autoregressive models for density time series. Shang & Haberman (2020, 2025a, 2025b) considered the center log-ratio transformation,
While any PDF $f$ may be thought of as random element of a Hilbert space, it does not reside in linear space due to summability and non-negativity constraints (i.e. $f\geq 0$ and $\int f=1$); if $f$, $g$ are two PDFs and $a$, $b$ two real values, then the combination $a\,f\,+\,b\,g$ is not a PDF anymore. Consequently, the standard space of square-integrable functions $\mathcal{L}^{2}$ is not the most appropriate for representing PDFs. A natural way to deal with such constraints is to peel them away by an invertible transformation that maps densities onto a linear space. From an extrinsic viewpoint, such transformations include the Bayes Hilbert space approach (see, e.g., Hron et al., 2016; Menafoglio et al., 2022), compositional data analysis (CoDa) (Shang & Haberman, 2020), cumulative distribution function transformation (Shang & Haberman, 2025a), $\alpha$ transformation (Shang & Haberman, 2025b) and log quantile density transformation (LQDT) (see, e.g., Petersen & Müller, 2016; Han et al., 2020). Kokoszka et al. (2019) compared the finite-sample performance of density estimation and forecasting among the LQDT, CoDa, unconstrained functional principal component regression of Horta & Ziegelmann (2018), and a skewed-$t$ distribution of Wang (2012). Kokoszka et al. (2019) recommended the LQDT and CoDa, two benchmark methods in our empirical studies.
In this paper, we focus on a random process taking values in some space of densities. We aim to forecast nonparametrically a future density given those observed in the past. Density forecasting is of great importance, especially in demography. In the realm of actuarial statistics, an important use of density forecasting is to predict the age distribution of life-table death counts as a means of computing survival probabilities. In such a model, any wrong specification of the forecast density may produce an inaccurate estimate of value-at-risk and make the asset holder, i.e., insurance companies, unable to control risk (see, e.g., Shang, 2013). In Figure 1, we present a time series of age-specific life-table death counts for the United States of America (USA) from 1933 to 2023 obtained from the Human Mortality Database (2025).
Figure 1: Rainbow plots of the age distribution of life-table death count from 1933 to 2023 in a single-year group. The life-table radix is 100,000 for each year. The life-table death counts in the oldest years are shown in red, while the most recent years are in violet. Curves are ordered chronologically according to the colours of the rainbow.
This setting faces three non-standard frameworks: 1) nonparametric density-on-density regression, 2) the dependency of the sequence of observed density functions, and 3) the convexity property of the space of densities that do not form a linear space. To take into account the nonlinear feature of the space of PDFs, we consider the Bayes Hilbert space approach (see, e.g., Egozcue et al., 2006; Van den Boogaart et al., 2010, 2014; Hron et al., 2016). The Bayes Hilbert spaces provide a mathematical environment allowing the extension of the Aitchison geometry from finite-dimensional compositional data (Aitchison, 1986; Pawlowsky-Glahn et al., 2015) to an infinite-dimensional setting. The Bayes Hilbert space is a linear space of PDFs designed for statistical models involving random PDFs (see, e.g., Delicado, 2011; Hron et al., 2016; Seo & Beare, 2019; Joen & Park, 2020).
A nonparametric density-on-density regression, where the response is one step ahead of the predictor in a time series of densities, is defined in the Bayes Hilbert space. A distance-based approach with a kernel-type estimator is proposed. Its flexibility makes the implementation attractive as various distance metrics may be considered. Its simplicity also makes the implementation easy and computationally fast as it does in the more standard nonparametric function-on-function regression setting (see, e.g., Ferraty et al., 2012; Lian, 2012). The advantage of nonparametric modelling is to capture the possible nonlinear dependence between two consecutive random PDFs. Unfortunately, instead of observing the density-valued random process, we sometimes only have samples drawn from these density functions (see also Kokoszka et al., 2019; Shang & Haberman, 2020; Zhang et al., 2022). This is why our methodology often includes, in its first step, the estimation of the densities before estimating the nonparametric relationship between densities.
The paper is organised as follows. Section 2.1 reminds the basic geometrical concepts of Bayes Hilbert space. In Section  2.2, we describe a model, estimation and forecasting scheme for our nonparametric density-on-density regression. In Section 2.3, we present asymptotic properties associated with our Bayes Nadaraya-Watson (NW) estimator. In Section 3.1, we describe how we estimate density via a kernel density estimator, the selection of optimal bandwidth, and the criterion used to assess density estimation accuracy. In Section 3.2, we evaluate and compare our Bayes NW estimator in the nonparametric density-on-density regression via a series of simulation studies. A French COVID19 hospitalisation data analysis is shown in Section 3.3, and a data analysis of age-specific life-table death count in the United States (U.S.) is presented in Section 3.4. The conclusion is given in Section 4, along with some ideas on how the methodology can be further extended.
2 Nonparametric density-on-density regression
We first briefly remind the definition of the Bayes Hilbert space before introducing the density-on-density regression model. It is essential to understand how this vector space of PDFs works and its associated geometry.
2.1 Bayes Hilbert space: a vector space of PDFs
Bayes Hilbert space.
Let $\mathcal{B}^{2}(I)$ be the set of bounded PDFs $f$ with continuous support $I=[a,\,b]$ for which the natural logarithm is square integrable:
$\mathcal{B}^{2}(I)=\left\{f:I\mapsto]0,\,+\infty[,\,\int_{I}f=1,\,\int_{I}\{\ln(f)\}^{2}\,<\,+\infty\right\}.$
Remark: to simplify notations, when there is no ambiguity, arguments of integrand are forgotten. Two fundamental operators $\oplus$ and $\odot$ respectively called perturbation and powering are defined for any $f$, $g$ in $\mathcal{B}^{2}(I)$ and $r$ in $\mathbb{R}$:
[…]
The perturbation is commutative and associative:
$f\oplus g=g\oplus f,\qquad(f\oplus g)\oplus h=f\oplus(g\oplus h).$
Let ${\bf 0}_{\mathcal{B}}=\frac{1}{b-a}1_{I}$, the uniform PDF defined on $I$, is the neutral element ($f\oplus{\bf 0}_{\mathcal{B}}={\bf 0}_{\mathcal{B}}\oplus f=f$) for perturbation and $f^{-1}$ is the unique element such that $f\oplus f^{-1}=f^{-1}\oplus f={\bf 0}_{\mathcal{B}}$. The power is associative ($r\odot(s\odot f)=(r\,s)\odot f$), distributive with respect to perturbation ($r\odot(f\oplus g)=(r\odot f)\oplus(r\odot g))$ and scalar addition ($(r+s)\odot f=(r\odot f)\oplus(s\odot f)$), and 1 is its neutral element. Based on the perturbation and powering, a third operator $\ominus$ called perturbation-substraction is defined:
$f\ominus g=f\oplus(-1)\odot g.$
In $\mathcal{B}^{2}(I)$, the inner product is defined as:
$\langle f,g\rangle_{\mathcal{B}}=\frac{1}{2\,(b-a)}\int_{I}\int_{I}\ln\frac{f(u)}{f(v)}\,\ln\frac{g(u)}{g(v)}\,du\,dv$
and one has useful properties: 1) $(\mathcal{B}^{2}(I),\,\oplus,\,\odot,\,\langle\cdot,\cdot\rangle_{\mathcal{B}})$is a vector space, 2) $(\mathcal{B}^{2}(I),\,\oplus,\,\odot,\,\langle\cdot,\cdot\rangle_{\mathcal{B}})$forms a separable Hilbert space called Bayes Hilbert space (see Egozcue et al., 2006; Van den Boogaart et al., 2014). At last, for any $f$, $g$ in $\mathcal{B}^{2}(I)$, $\|f\|_{\mathcal{B}}=\sqrt{\langle f,f\rangle_{\mathcal{B}}}$ is the norm associated with the inner product. One can remark that $\|f\ominus g\|_{\mathcal{B}}=0$ is equivalent to $f\ominus g={\bf 0}_{\mathcal{B}}$, which results in $f(u)=g(u)$ for any $u\in I$.
Isometry between $\mathcal{B}^{2}(I)$ and $\mathcal{L}^{2}_{0}(I)$.
Let $\mathcal{L}^{2}_{0}(I)$ be the space of square-integrable real functions $f$ on $I$ with $\int_{I}f=0$. There exists an isometric isomorphism between $\mathcal{B}^{2}(I)$ and $\mathcal{L}^{2}_{0}(I)$ called centered $\log$-ratio (clr) transformation, defined for any PDF $f$ in $\mathcal{B}^{2}(I)$ and any $u$ in $I$:
$\text{clr}(f)(u)=\ln f(u)-(b-a)^{-1}\int_{I}\ln f.$
In addition of the isometry property of the clr transformation $\|f\ominus g\|_{\mathcal{B}}=\|\text{clr}(f)-\text{clr}(g)\|$ where $\|g\|=\int_{I}g^{2}$, $\text{clr}(f\oplus g)=\text{clr}(f)+\text{clr}(g)$, $\text{clr}(s\odot f)=s\times\text{clr}(f)$, and its inverse transformation
$\text{clr}^{-1}(g)=\frac{\exp(g)}{\int_{I}\exp(g)},$
for any $g$ in $\mathcal{L}^{2}_{0}(I)$.
Once the Bayes Hilbert space framework is reminded, one can introduce our nonparametric density-on-density regression when considering a time series of PDFs.
2.2 Model, estimation, and sequential forecasting scheme
2.2.1 Density-on-density regression model
Let $\{\mathscr{F}_{t}\}_{t\in\mathbb{Z}}$ be a stochastic process taking values in $\mathcal{B}^{2}(I)$ and consider $N+1$ random PDFs $(f_{1},\ldots,f_{N+1})$ as realizations of $(\mathscr{F}_{1},\ldots,\mathscr{F}_{N+1})$. The main goal is to forecast a future PDF $f_{N+2}$ given those observed in the past ($f_{N+1},\,f_{N},\ldots,f_{1}$). To this end, we propose the nonparametric density-on-density regression model
$f_{t+1}=m(f_{t})\oplus\epsilon_{t},\qquad t=1,\dots,N,$ (1)
where $m(\cdot)$ is an unknown smooth operator mapping $\mathcal{B}^{2}(I)$ into $\mathcal{B}^{2}(I)$, and for each $t$, $\text{E}(\epsilon_{t}|f_{t})={\bf 0}_{\mathcal{B}}$, where the conditional expectation is defined for $\mathcal{B}^{2}(I)$-valued random PDFs. The definition of conditional expectation for Hilbert-valued random variable can be found in Bosq (2000). Under Model (1), it is easy to check that, for each $t$, $\text{E}(\epsilon_{t}|f_{t})={\bf 0}_{\mathcal{B}}$ is equivalent to $m(f)=\text{E}\left(f_{t+1}|f_{t}=f\right)$ (see Section A in Appendix). Our main challenge is to define for any given density $f$ an estimator $\widetilde{m}(f)$ of the unknown regression operator $m(f)=\text{E}\left(f_{t+1}|f_{t}=f\right)$ fulfilling density features (nonnegative functions that integrate to one) and good asymptotic behaviour while no reducing functional form (i.e. linearity, additivity, functional index model, etc) of the unknown regression operator $m(\cdot)$ is assumed.
2.2.2 Kernel estimator
With some smoothness properties, the functional form of $m(\cdot)$ is often estimated in a data-driven manner. There are a number of nonparametric functional estimators, such as functional Nadaraya-Watson (NW) estimator (Ferraty & Vieu, 2006), functional local linear estimator (Berlinet et al., 2011; Ferraty & Nagy, 2022), functional $k$-nearest neighbour estimator (Burba et al., 2009, for scalar response and Lian, 2011, for functional response), functional smoothing splines (Crainiceanu & Goldsmith, 2010), functional wavelet estimator (Antoniadis & Sapatinas, 2003), distance-based local linear estimator (Boj et al., 2010) and functional neural network (Rossi & Villa, 2006). Throughout the paper, we introduce the functional Bayes NW estimator, i.e., the functional NW estimator in the Bayes-Hilbert space, because of its simplicity and mathematical elegance. The essence of functional Bayes NW kernel smoothing is to allow flexible estimation of the unknown regression operator. The functional Bayes NW estimator of the conditional mean can be defined as
$\widetilde{m}_{N}(f)\,=\,\bigoplus^{N}_{t=1}w_{h_{r}}(f_{t},\,f)\,\odot\,f_{t+1}\ \mbox{ with }\ w_{h_{r}}(f_{t},\,f)\,=\,\frac{K_{r}(h_{r}^{-1}\,\|f_{t}\ominus f\|_{\mathcal{B}})}{\sum^{N}_{t=1}K_{r}(h_{r}^{-1}\,\|f_{t}\ominus f\|_{\mathcal{B}})},$ (2)
where $K_{r}(\cdot)$ is an asymmetric kernel function and $h_{r}$ is a positive smoothing parameter (i.e. bandwidth). The vector space property of $\mathcal{B}^{2}(I)$ ensures that the Bayes NW estimator is a PDF. Depending on the kernel function and bandwidth parameter, the nonparametric estimator in (2) is a weighted average of the observed densities and has two interesting properties: it fulfils the PDF constraints by construction, and due to its simplicity, its implementation is straightforward and computationally fast.
In practice, conditionally on the densities $f_{1},\ldots,f_{N+1}$, we observe $n_{t}$ copies $X_{t,1},\ldots,X_{t,n_{t}}$ that share the same marginal density $f_{t}$ for each $t=1,\ldots,N+1$. In other words, the PDFs are not directly observable. This situation is particularly common in epidemiology. For instance, suppose that at each date $t$, COVID-19 confirmed cases corresponding to $\omega$ French departments are observed. In this case, $X_{t,j}$ is the COVID-19 confirmed cases at date $t$ of the department $j$ and all the sample sizes are equal $(n_{1}=n_{2}=\dots=\omega)$. This is why our methodology first focuses on kernel estimators $\widehat{f}_{1},\ldots,\widehat{f}_{N+1}$ of the $N+1$ probability densities $f_{1},\ldots,f_{N+1}$. In the second step, we just plug the estimated densities $\widehat{f}_{1},\ldots,\widehat{f}_{N+1}$ into the nonparametric density-on-density regression estimator (2) which provides our definitive kernel estimation:
$\widehat{m}_{N}(f)\,=\,\bigoplus^{N}_{t=1}w_{h_{r}}(\widehat{f}_{t},\,f)\,\odot\,\widehat{f}_{t+1},$
where $w_{h}(\widehat{f}_{t},\,f)$ is defined in (2). A standard by-product of this estimating procedure in the time series context is the following one-step-ahead sequential forecasting scheme. From the estimated density functions $\widehat{f}_{1},\ldots,\widehat{f}_{N+1}$, we compute $\widehat{f}_{N+2}=\widehat{m}_{N}(\widehat{f}_{N+1})$ which is the prediction of the density $f_{N+2}$. To forecast the next density, just consider the sample of random densities completed with the predicted one (i.e. $\widehat{f}_{1},\ldots,\widehat{f}_{N+1},\widehat{f}_{N+2}$) and build $\widehat{f}_{N+3}=\widehat{m}_{N+1}(\widehat{f}_{N+2})$. This one-step-ahead forecasting scheme can be iterated to get $\widehat{f}_{N+\mathcal{H}}$ at any reasonable forecast horizon $\mathcal{H}$.
2.2.3 Dependence model
To better understand the dependence structure of our data, our framework combines two random mechanisms: 1) the underlying process {}$\{\mathscr{F}\}_{t\in\mathbb{Z}}$ generating the random PDFs $\{f_{1},\ldots,f_{N+1}\}$ and 2) the random process providing the $N+1$ samples $\left\{X_{t,1},\ldots,X_{t,n_{t}}\right\}_{t=1,\ldots,N+1}$ for estimating the non-observable random PDFs. For any $t=1,2,\ldots,N$, set $\bm{X}_{t}=\left\{X_{t,1},\ldots,X_{t,n_{t}}\right\}$; the scheme hereafter sums up the overall random situation:
Density-valued process $f_{1}$longrightarrow $f_{2}=m(f_{1})\oplus\epsilon_{1}$longrightarrow $\cdots$ $\longrightarrow$ $f_{N+1}=m(f_{N})\oplus\epsilon_{N}$
[…]
Observed data $\bm{X}_{1}$ $\bm{X}_{2}$ $\bm{X}_{N+1}$
[…]
Estimated densities $\widehat{f}_{1}$ $\widehat{f}_{2}=\widehat{m}_{N}(\widehat{f}_{1})$ $\widehat{f}_{N+1}=\widehat{m}_{N}(\widehat{f}_{N})$
We assume that the random process of two consecutive densities $\{(\mathscr{F}_{t},\mathscr{F}_{t+1})\}_{t\in\mathbb{Z}}$ is a strictly stationary $\rho$-mixing process (see, e.g., Rio, 1993; Bradley, 2005). We first remind the definition of the $\rho$-mixing dependence condition. For any $\sigma$-field $\mathcal{A}$, let $\mathcal{L}^{2}(\mathcal{A})$ denote the space of square-integrable $\mathcal{A}$-measurable real-valued random variables. Given two $\sigma$-fields $\mathcal{A}$ and $\mathcal{B}$, $\rho(\mathcal{A},\mathcal{B})=\sup\big{\{}\left|\text{Corr}(U,\,V)\right|$, $\,U\in\mathcal{L}^{2}(\mathcal{A}),\,V\in\mathcal{L}^{2}(\mathcal{B})\big{\}}$ measures the dependence between $\mathcal{A}$ and $\mathcal{B}$ in terms of correlation. $\rho$ is between 0 and 1; the closer this quantity is to 0, the weaker the dependency. Let $\{\mathscr{Z}_{t}\}_{t\in\mathbb{Z}}$ be a random process and for any $k<\ell$, define $\mathcal{A}_{k}^{\ell}$ as the $\sigma$-filed generated by the $\mathscr{Z}_{t}$’s for $t$ between $k$ and $\ell$. The random sequence $\{\mathscr{Z}_{t}\}_{t\in\mathbb{Z}}$ is said to be $\rho$-mixing if the dependence coefficients $\rho(m)=\sup_{j\in\mathbb{Z}}\rho\left(\mathcal{A}_{-\infty}^{j},\,\mathcal{A}_{j+m}^{+\infty}\right)$ tends to 0 as $m$ goes to infinity. This asymptotic property is usually called the $\rho$-mixing condition. Assuming that $\{(\mathscr{F}_{t},\mathscr{F}_{t+1})\}_{t\in\mathbb{Z}}$ is $\rho$-mixing is equivalent to suppose that its dependence coefficients fulfill the $\rho$-mixing condition with $\mathscr{Z}_{t}=(\mathscr{F}_{t},\mathscr{F}_{t+1})$. Then, according to the scheme generating the data, it is reasonable to suppose that the random sequences $(f_{t},\,f_{t+1})_{t=1,2,\dots}$ and $(\bm{X}_{t},\,\bm{X}_{t+1})_{t=1,2,\dots}$ inherit the structure (i.e strictly stationary and $\rho$-mixing) of $\{(\mathscr{F}_{t},\mathscr{F}_{t+1})\}_{t\in\mathbb{Z}}$. As a by-product, the random sequence $(\widehat{f}_{t},\,\widehat{f}_{t+1})_{t=1,2,\dots}$ based on the estimated densities is also a strictly stationary $\rho$-mixing process. Lastly, to complete the structure of our data, conditionally to $f_{t}$ and for each $t=1,\ldots,N$, the error model $\epsilon_{t}$ is assumed independent of $\bm{X}_{t}$.
2.3 Asymptotic properties
This section is devoted to the theoretical properties of our density’s functional kernel estimator. We first introduce notations and assumptions we need to derive the asymptotic behaviour of $\widehat{m}(f)$.
  • (H1)
    It exists $C>0$ such that, for all $f$ and $g$ in $\mathcal{B}^{2}(I)$, $\|m(f)\ominus m(g)\|_{\mathcal{B}}\,\leq\,C\,\|f\ominus g\|_{\mathcal{B}}$,
  • (H3)
    it exists $a>1$ such that the $\rho$-mixing coefficient fulfills the condition $\rho(k)=O(k^{-a})$,
  • (H5)
    set $n=\inf_{t\geq 1}n_{t}$ and let $n=n_{N}$ be a sequence depending on $N$ such that $n$ goes to infinity with $N$; one assumes the existence of a sequence $\delta_{N}$ tending to 0 as $N$ goes to infinity such that, for any $t$, $\text{E}\left(\|\widehat{f}_{t}\ominus f_{t}\|^{2}_{\mathcal{B}}\right)=O(\delta_{N}^{2})$,
  • (H7)
    For any deterministic $f$ in $\mathcal{B}^{2}(I)$, set $\pi_{f}(h_{r})=P\left(\|f_{1}\ominus f\|_{\mathcal{B}}<h_{r}\right)$; it exists $C>0$ and $C^{\prime}>0$, $0<C\pi_{f}(h_{r})\leq\pi_{f}\left(h_{r}+o(h_{r})\right)\leq C^{\prime}\pi_{f}(h_{r})$,
  • (H9)
    $h_{r}$ tends to 0 with $N$, $N^{a/(1+a)}\,\pi_{f}(h_{r})$ goes to infinity with $N$, and it exists $b$ with $11/3<b<$ 1 such that $\delta_{N}^{2b}=o\left(\pi_{f}(h_{r})\right)$ and $\delta_{N}^{1-b}=o\left(h_{r}\right)$,
  • (H11)
[…]
(H(H1)) is usual; it requires that the regression operator is smooth enough by fulfilling the Lipschitz property. In (H(H3)), the model dependency assumes an arithmetic decay for the mixing coefficients. This assumption simplifies the convergence rate, but more general writing involving the mixing coefficients $\rho(N)$ can be obtained. (H(H5)) relates, for each $t$, the size $n_{t}$ of the sample $X_{t,1},\ldots,X_{t,n_{t}}$ to the number $N$ of random PDFs. In addition, it requires that the mean integrated squared error $\text{E}\left(\|\widehat{f}_{t}\ominus f_{t}\|_{\mathcal{B}}^{2}\right)$ of $\widehat{f}_{t}$ tends to zero when $n_{t}$ (i.e. $N$) goes to infinity. In the standard multivariate setting, the sequence $\delta_{N}$ can be interpreted as an upper bound of the well-known quantity $\text{MISE}(\widehat{f}_{t})=\text{E}\left(\|\widehat{f}_{t}-f_{t}\|\right)$ (for more details on MISE, see Rosenblatt, 1956; Epanechnikov, 1969; Wegman, 1972; Nadaraya, 1974; Marron & Wand, 1992; Hansen, 2005, among others). It is shown Section A in Appendix that as soon as $f\in\mathcal{B}^{2}(I)$ with $f\geq\theta>0$, $\text{E}\left(\|\widehat{f}_{t}\ominus f_{t}\|_{\mathcal{B}}^{2}\right)=O\left\{\text{MISE}(\widehat{f}_{t})\right\}$. In other words, when the density $f$ admits a strictly nonnegative lower bound, then $\text{E}\left(\|\widehat{f}_{t}\ominus f_{t}\|_{\mathcal{B}}^{2}\right)$ inherits the asymptotic properties of MISE. The hypothesis (H(H7)) assumes that the behaviour of the small ball probability $\pi_{f}(h)$ is not sensitive to variations on radius negligible with respect to the bandwidth $h$ (see Li & Shao, 2001 for a survey on small ball probability). According to the isometry property of the clr transformation (see Section 2.1), $\pi_{f}(h_{r})=P(\|X-x\|_{2}<h_{r})$ with $X=\text{clr}(f_{1})$ and $x=\text{clr}(f)$ in $\mathcal{L}^{2}_{0}(I)$; the class of $\mathcal{L}^{2}_{0}(I)$-valued random functions fulfilling (H(H7)) encompasses Gaussian-type random processes (see Ferraty et al., 2019). (H(H9)) relates the asymptotic behaviour of $h_{r}$, $\pi_{f}(h_{r})$, $N$ and $\delta_{N}^{2}$, the upper bound of $\text{E}\left(\|\widehat{f}_{t}\ominus f_{t}\|_{\mathcal{B}}^{2}\right)$ for any $t$. As $\pi_{f}(h_{r})$ tends to zero with $N$ faster than $h_{r}$ does, it is reasonable to require that $02b>1-b>$ 0 which results in $11/3<b<$ 1. With (H(H11)) and for any $q>0$, $\text{E}\left\{K^{q}_{r}(h^{-1}\,\|f_{t}\ominus f\|_{\mathcal{B}})\right\}$ is bounded below and above by a quantity of order $\pi_{f}(h_{r})$. For simplicity, we use a basic hypothesis on the kernel function. Nevertheless, it is possible to get the more accurate result $\text{E}\left\{K^{q}_{r}(h_{r}^{-1}\,\|f_{t}\ominus f\|_{\mathcal{B}})\right\}\,=\,C_{f,q}\pi_{f}(h_{r})\{1+o(h_{r})\}$ as soon as the kernel function $K_{r}$ is continuously differentiable on $(0,1)$ with $K_{r}(1)>0$, and for all $s\in(0,1)$, $K_{r}^{\prime}(s)\leq 0$ and the ratio $\pi_{f}(h_{r}\,s)/\pi_{f}(h_{r})$ tends to some positive value as $h$ tends to 0.
Once notations and assumptions are introduced, one can provide our main result by giving the pointwise convergence rate of our density-on-density kernel regression estimator; the proof is postponed to Section B of the Appendix.
Theorem 1
If (H(H1))-(H(H11)) are fulfilled, then
$\left\|\widehat{m}_{N}(f)\ominus m(f)\right\|_{\mathcal{B}}\,=\,O(h_{r})+O\left(\delta_{N}\,\pi_{f}(h_{r})^{-1/2}\right)+O_{P}\left(\left\{N^{a/(1+a)}\,\pi_{f}(h_{r})\right\}^{-1/2}\right).$
The first two terms correspond to the bias part. $O(h_{r})$, which is not surprising (see Ferraty & Vieu, 2006), is derived from the regularity assumption acting on the regression operator $m$ (see (H(H1))). The second term, which is not usual, involves $\delta_{N}$ and the small ball probability $\pi_{f}(h_{r})$. This is the price to pay when the random PDFs are not directly observed and are approximated by means of some estimating procedure, where $\delta_{N}$ can be interpreted as the upper bound of the root MISE of $\widehat{f}_{t}$ for any $t$. For the variance part, remember that $a$ controls the asymptotic behaviour of the mixing coefficient. So, the impact of the dependency model appears directly in the exponent of the sample size $N$ of the random densities. In the situation of a sample of independent explanatory functional random variables, we would get $N$ instead of $N^{a/(1+a)}$ (see again Ferraty & Vieu, 2006). As expected, the dependence structure of the data degrades the rate of convergence of $\widehat{m}$.
Let us now focus on the estimation of the random density $f_{t}$ with a particular attention on the sequence $\delta_{N}$ which is an upper bound of $\text{MISE}(\widehat{f}_{t})$. The more standard situation corresponds to the observation of an independent and identically distributed (iid) sample $X_{t,1},\ldots,X_{t,n_{t}}$ but dependent data received also lots of attention (see Rosenblatt, 1971; Nadaraya, 1974; Chanda, 1983; Hart, 1984; Castellana & Leadbetter, 1986; Robinson, 1986; Hall & Hart, 1990; Chesneau, 2014, among others). In the special case of a kernel density estimator
$\widehat{f}_{t}(x)=(n_{t}\,h_{t})^{-1}\sum_{i=1}^{n}K_{d}\left\{h_{t}^{-1}(x-X_{t,i})\right\},$
with $h_{t}$ a bandwidth tending to zero when $n_{t}$ goes to infinity and $K_{d}$ a kernel function, the asymptotic behaviour of $\text{MISE}(\widehat{f}_{t})$ is well known, even when the data $X_{t,1},\ldots,X_{t,n_{t}}$ come from a stationary strong mixing process. When $f_{t}$ is deterministic, Biau (2002) stated that the asymptotic behaviour of $\text{MISE}(\widehat{f}_{t})$ for dependent observations (with arithmetic decay rate for the mixing coefficient) is the same than those obtained for independent data: $\text{MISE}(\widehat{f}_{t})\,=\,O\left(h_{t}^{4}\right)+O\left(n_{t}^{-1}h_{t}^{-1}\right)$. Setting $h_{t}\sim n_{t}^{-1/5}$ in order to balance both terms results in $\text{MISE}(\widehat{f}_{t})\,=\,O\left(n_{t}^{-4/5}\right)$. This result requires that the first two derivatives $f_{t}^{\prime}$ and $f_{t}"$ of the true density function $f_{t}$ are continuous with $\int({f_{t}"})^{2}<\infty$. When $f_{t}$ is a random density, one can write $\text{MISE}(\widehat{f}_{t})\,=\,\text{E}\left\{\text{E}\left(\|\widehat{f}_{t}-f_{t}\|^{2}|f_{t}\right)\right\}$. Then, it is easy to see that the same result holds as soon as
  • (H13)
    $f_{1}^{\prime}$ and $f_{1}"$ are continuous almost surely with $\text{E}\left\{\int\left(f_{1}"\right)^{2}\right\}<\infty$.
The following result simplifies the rate of convergence given in Theorem 1 when the random densities are estimated with the kernel density estimator.
Corollary 1
If, in addition of (H(H1))-(H(H13)), one has
  • (H15)
    $X_{t,1},\ldots,X_{t,n_{t}}$ are iid or generated from a stationary $\rho$-mixing process with arithmetic decay rate for the mixing coefficients,
  • (H17)
    for any $t$, it exists $q>5a/\{4b(1+a)\}$ such that $n_{t}\sim N^{q}$ and $h_{t}\sim N^{-q/5}$,
Then
$\left\|\widehat{m}_{N}(f)\ominus m(f)\right\|_{\mathcal{B}}\,=\,O(h_{r})+O_{P}\left(\left\{N^{a/(1+a)}\,\pi_{f}(h_{r})\right\}^{-1/2}\right).$
(H(H17)) corresponds to a particular case where $\delta_{N}^{2}$, the upper bound of $\text{E}\left(\|\widehat{f}_{t}\ominus f_{t}\|_{\mathcal{B}}^{2}\right)$, is negligible with respect to $N^{-a/(1+a)}$ so that the term $O\left(\delta_{N}\,\pi_{f}(h_{r})^{-1/2}\right)$ vanishes.
3 Finite-sample properties
3.1 Implementation
Since the actual densities in practical applications are not often observable, we work with densities which are outputs of kernel density estimators:
$f_{t}(u)=\frac{1}{n_{t}h_{t}}\sum^{n_{t}}_{i=1}K\left(\frac{u-X_{t,i}}{h_{t}}\right),\qquad t=1,\dots,n,$ (3)
where $K(\cdot)$ denotes a kernel function and $h_{t}$ denotes a bandwidth for period $t$. We consider the truncated Gaussian kernel.
In our nonparametric density-on-density regression, we consider how to select the optimal bandwidths for estimating the nonparametric regression operator in (1) and for estimating densities in (3). To estimate densities from observed data, we use a kernel density estimator with the bandwidth selected by Silverman’s rule-of-thumb (ROT) (see also Kokoszka et al., 2019). Kokoszka et al. (2019) used Silverman’s ROT to select the bandwidth, which leads to $\widehat{h}_{t}=2.34\times\widehat{\sigma}_{t}\times n_{t}^{-1/5}$, where $\widehat{\sigma_{t}}$ denotes the sample standard deviation of the returns $X_{t,i}$, $i=1,\dots,n_{t}$.
To estimate the bandwidth for the nonparametric regression operator, we can use generalised cross-validation (GCV) to select the bandwidth automatically (see, e.g., Ferraty et al., 2012). The GCV aims to minimise an overall squared loss function between the estimated and observed responses among the data in the training sample.
3.2 Monte-Carlo simulation studies
We conduct a series of simulation studies to examine the finite-sample performance of the nonparametric density-on-density regression described in Section 2. Our data generating process consists of the following steps:
  1. 1)
    When generating the model $f_{t+1}=m(f_{t})\oplus\epsilon_{t}$, a first step is to be able to simulate $\epsilon_{0},\epsilon_{1},\ldots$, the probability density functions playing the role of model errors. By using the clr transformation, the density-on-density regression model is equivalent to $\mbox{clr}(f_{t+1})=\mbox{clr}\circ m(f_{t})+\mbox{clr}(\epsilon_{t})$. Instead of generating directly the errors $\epsilon_{t}$ in the Bayes Hilbert space $\mathcal{B}^{2}(I)$, it could be more convenient to simulated functional errors $\eta_{t}=\mbox{clr}(\epsilon_{t})$ in the corresponding Hilbert space $\mathcal{L}^{2}_{0}(I)$.
  2. 2)
[…]
    $\eta_{t}(u)\,=\,\sum_{k\in\mathcal{K}}A_{tk}\,\phi_{k}(u).$
[…]
    $\eta_{t}(u)\,=\,A_{t1}\cos(\pi\,u)+A_{t2}\sin(\pi\,u)+A_{t3}\cos(2\pi u)+A_{t4}\sin(2\pi\,u)+A_{t5}\cos(3\pi\,u),$
    where $A_{tj}$’s are iid zero-mean real random variables (r.r.v.). In Figure 2, we present such model errors $\eta_{t}$ with $A_{t1}$’s,…,$A_{t5}$’s $\overset{iid}{\sim}$ $N(0,\sigma^{2})$.
    Figure 2: The basis functions we considered are the first five trigonometric functions. By multiplying them with random coefficients simulated from a Gaussian distribution, we obtain the model error in the clr space, $\eta_{t}(u)$.
    Transform back to $\epsilon_{t}$’s with the inverse of the clr transformation: $\epsilon_{t}=\mbox{clr}^{-1}(\eta_{t})$. In Figure 3, we present 10 first model errors in Bayes space.
    Figure 3: By taking the inverse clr transformation of the model error in Figure 2, we obtain the model error in Bayes space.
  3. 3)
    Let us define the regression operator $m:\,\mathcal{B}^{2}(I)\rightarrow\mathcal{B}^{2}(I)$ such that
[…]
    with $f$ and $g$ two given probability densities in $\mathcal{B}^{2}(I)$, $10<\rho_{0}<$ 1 and where $\underline{b}=\max[-1$, $\rho_{0}^{-1}(y+\rho_{0}-1)]$ and $\overline{b}=\min[1,\rho_{0}^{-1}(y-\rho_{0}+1)]$. Depending on the value of $\rho_{0}$, the function support may vary for the regression mean function. This regression operator has a nice interpretation in terms of convolution. Suppose that, for any $10<\rho_{0}<$ 1, $Z=\rho_{0}X+(1-\rho_{0})Y$ with $X$ and $Y$ two independent r.r.v. and let $f_{Z}$ (resp. $f_{X}$) be the PDF of $Z$ (resp. $X$). Then, we have $f_{Z}\,=\,m(f_{X})$ and $\text{Corr}(Z,\,X)=\rho_{0}$. The regression operator $m(.)$ corresponds to the transformation resulting from a convex combination of $X$ with another r.r.v. $Y$. Based on this remark, we propose the following scheme for simulating our time series of PDFs in the Bayes Hilbert space:
    $f_{t+1}\,=m(f_{t})\oplus\epsilon_{t},$
    where the $\epsilon_{t}$’s are the model errors in Bayes space (built previously): 1) $m(f_{t})$represents the PDF of $X_{t+1}=\rho_{0}X_{t}+(1-\rho_{0})Y_{t+1}$ when $f_{t}$ is the PDF of $X_{t}$, 2) $f_{t+1}$is the obtained PDF when adding some model error $\epsilon_{t}$ to $m(f_{t})$.
  4. 4)
    Set $Y_{t}\,\sim\,g_{t}=TN(\mu_{t},\nu^{2})$ where $TN(\mu_{t},\nu^{2})$ is the truncated normal distribution over $[-1,1]$ derived from the normal distribution $N(\mu_{t},\nu^{2})$ where, for some constant $T>0$, $\mu_{t}=\cos(2\pi t/T)$.
  5. 5)
    Set $f_{1}\equiv g_{1}$ where $g_{1}=TN(\mu_{1},\nu^{2})$ with $\mu_{1}=\cos(2\pi/T)$. We compute $f_{2}=m(f_{1})\oplus\epsilon_{1}$. Then, we iterate to build the whole time series of densities $f_{1},\,f_{2}=m(f_{1})\oplus\epsilon_{1},\ldots,\,f_{N}=m(f_{N-1})\oplus\epsilon_{N-1}$.
In Figure 4, we present a perspective plot of a simulated example with $150$ curves. We consider 201 grid points within a function support range $[-1,1]$. There are four tuning parameters to choose, namely $\sigma$ in step 2), $\rho_{0}$ in step 3), $\nu$ and $T$ in step 4). Let $\sigma=0.1$, $\rho_{0}=0.5$, $\nu=0.5$ and $T=150$.
Figure 4: 150 simulated densities with 201 grid points bounded between -1 and 1. The four tuning parameters are chosen as: $\sigma=0.1$, $\rho=0.5$, $\nu=0.5$ and $T=150$.
We divide the simulated data into a training sample and a testing sample comprised of the last 50 densities. To evaluate the finite-sample performance, we compute the Kullback-Leibler divergence (KLD) (Kullback & Leibler, 1951). The KLD is intended to measure the loss of information when we choose an approximation. For two density functions, the discrete version of the Kullback-Leibler divergence is defined as
$\text{KLD}_{m}$ $=\frac{1}{50}\sum^{50}_{i=1}\left\{D_{\text{KL}}[f_{i}^{m}(s)||\widehat{f}_{i}^{m}(s)]+D_{\text{KL}}[\widehat{f}_{i}^{m}(s)||f_{i}^{m}(s)]\right\}$
$\overline{\text{KLD}}$ $=\frac{1}{100}\sum^{100}_{m=1}\text{KLD}_{m},$
where $f_{i}^{m}(s)$ is the $i$th observation in the testing sample, and $\widehat{f}_{i}^{m}(s)$ is the one-step-ahead forecast of $f_{i}^{m}(s)$ for the $m$th replication.
For one out of 100 replications, we generate a sample of densities under different tuning parameters. With the training samples, we produce one-step-ahead forecast densities using the nonparametric density-on-density regression with the Bayes NW estimator, the CoDa method, the LQDT, and functional principal component regression of Horta & Ziegelmann (2018). In Figure 5, we present one replication of holdout densities and their forecasts obtained by the five methods. For this example, we observe that the nonparametric density-on-density regression produces the best results, followed closely by the random walk. The Horta & Ziegelmann’s (2018) method suffers from the well-known tail problem when density forecasts can even be negative.
Figure 5: One simulated example with a sample size of $T=150$ and 201 grid points between -1 and 1. While the first 100 densities are used as the initial training sample, we produce iterative one-step-ahead forecast densities via an expanding window scheme. HZ stands for Horta & Ziegelmann’s (2018) method.
To further assess the overall forecast accuracy, we repeat the simulation data generating process 100 times, each with different pseudo-random seeds. In Table 1, we present the one-step-ahead averaged KLDs between the actual holdout samples and their corresponding forecasts obtained from the five methods. Further, we present results using expanding window approach where the size of the training samples increases. As measured by the $\overline{\text{KLD}}$ criterion, the forecast accuracy relies heavily on the noise-to-signal ratio, characterised by the dependence parameter $\rho_{0}$ and the standard deviation $\sigma$ in the error term of the AR(1) structure. As the dependence parameter $\rho_{0}$ decreases from 0.75 to 0.25, we observe a decrease of $\overline{\text{KLD}}$ based on 100 replications. As the standard deviation in the error term of the AR(1) structure increases from 0.1 to 1, we also observe an increase of $\overline{\text{KLD}}$. As the sample size increases from $150$ to 550, we observe a general increase of $\overline{\text{nsr}}$ which does not translate to an increase of $\overline{\text{KLD}}$.
Table 1: One-step-ahead forecast accuracy of the nonparametric density-on-density regression with the Bayes NW estimator in the simulated data with 100 replications. We also compare the forecast accuracy with the CoDa, LQDT, HZ and RW methods, under three choices of $(\sigma,\rho_{0})$ with sample sizes of $T=150,350$ and $550$. The last 50 densities are the testing sample, while the remaining densities are the training sample. For the 100 replications, we compute the average noise-to-signal ratio, denoted by $\overline{\text{nsr}}$.
Method
$(\sigma,\rho_{0})$ $n$ $\overline{\text{nsr}}$ Bayes NW CoDa LQDT HZ RW
$(0.10,0.75)$ 150 0.0398 0.0137 1.2085 3.7018 0.6377
0.0115
350 0.0918
0.0079
0.4419 3.5817 0.0818 0.0108
550 0.1439 0.0209 0.1213 0.0173 0.1772
0.0197
$(0.10,0.50)$ 150 0.0219
0.0069
0.1339 0.0880 0.0963 0.0097
350 0.0353
0.0035
0.0593 0.0155 0.0389 0.0094
550 0.0514
0.0082
0.0108 0.0119 0.0754 0.0125
$(0.10,0.25)$ 150 0.0272
0.0069
0.0170 0.2433 0.0628 0.0098
350 0.0340
0.0016
0.0079 0.1584 0.0516 0.0098
550 0.0434 0.0046
0.0032
0.0880 0.0501 0.0122
$(0.50,0.75)$ 150 0.9703
0.1638
1.1957 3.1611 0.6668 0.2371
350 2.2291
0.1200
0.4611 3.3656 0.3557 0.2334
550 3.4945
0.1376
0.1636 0.1967 0.4982 0.2717
$(0.50,0.50)$ 150 0.5338
0.0784
0.1924 0.1072 0.2682 0.2124
350 0.8582
0.0434
0.0901 0.1528 0.2076 0.2134
550 1.2480 0.0540
0.0396
0.1332 0.2796 0.2423
$(0.50,0.25)$ 150 0.6640 0.0522
0.0382
0.0994 0.2334 0.2191
350 0.8259 0.0233
0.0207
0.1421 0.1564 0.2189
550 1.0549 0.0304
0.0184
0.1772 0.1880 0.2499
$(1,0.75)$ 150 3.1793
0.5017
0.6227 3.3689 0.9971 0.7988
350 7.3052 0.3878
0.3745
3.3869 0.7850 0.8163
550 11.4313 0.3754
0.3416
0.5763 0.9343 0.8222
$(1,0.50)$ 150 1.7491
0.2599
0.3605 0.7276 0.4833 0.7203
350 2.8124
0.1561
0.1923 0.8041 0.4716 0.7303
550 4.0825 0.1743
0.1165
0.4573 0.5774 0.7667
$(1,0.25)$ 150 2.1756 0.1720
0.1144
0.7714 0.3197 0.7273
350 2.7067 0.0774
0.0615
0.8935 0.3034 0.7255
550 3.4510 0.1134
0.0572
0.5735 0.3942 0.7929
We observe that the Horta & Ziegelmann’s (2018) method produces inferior accuracy because it ignores the density constraints. By obeying the density constraints, the remaining four methods improve the forecast accuracy. Among the four methods, the LQDT does not work well when functions have different finite supports, as evident from large $\overline{\text{KLD}}$ values. When the $\sigma$ value is smaller, the nonparametric density-on-density regression using the functional Bayes NW estimator is the chosen method with the smallest $\overline{\text{KLD}}$. When the $\sigma$ value is larger, and $\rho_{0}$ value is smaller, the CoDa method is preferred.
3.3 Analysis of a French COVID19 hospitalisation dataset
The French open data portal https://www.data.gouv.fr/fr/ allows the general public to access many data sources. Simple research on this site with the keyword COVID19 leads us to several data sets. In this subsection, we are interested in the dynamics of the pandemic in France at the departmental level and, more particularly, in the hospitalisation data. French territory is divided into 101 administrative areas named départements. In Figure 6, we plot 96 departments located inside Europe.
Figure 6: A map of French departments
The French COVID19 data, updated daily, are contained in a .csv file whose name starts with covid-hospit-2023 and can be downloaded https://www.data.gouv.fr/fr/datasets/donnees-hospitalieres-relatives-a-lepidemie-de-covid-19/. We are interested in studying the number of hospitalisations per 100,000 inhabitants for each department from March 18, 2020, to August 24, 2022. By focusing on cross-sectional data, we collect 96 daily hospitalisation ratios $(\mathcal{X}_{t,1},\mathcal{X}_{t,2},\dots,\mathcal{X}_{t,96})$ for each observed date $t=1,2,\dots,890$. By taking a simple average over days, we can single out the departments with minimum and maximum hospitalisation ratios. The department Vendée has the minimum hospitalisation ratio, with an average of 8.58 per 100,000 inhabitants. The department Territoire de Belfort has the maximum hospitalisation ratio, with an average of 50.52 per 100,000 inhabitants.
Since the number of hospitalisation is nonnegative, it is an unnecessary constraint that can be removed via a natural logarithm transformation, denoted by $\ln(\cdot)$. Let $Z_{t,i}=\ln(\mathcal{X}_{t,i}+c)$ with $c>0$ for $i=1,2,\dots,96$, where $c$ is an arbitrary small value such as $c=0.1$. Then, we apply a kernel density estimator to estimate the probability density function $f_{t}$ of $Z_{t}$ without positive constraint. Computationally, the estimation is done via density function in [Uncaptioned image] (R Core Team, 2025) using Silverman’s ROT bandwidth selection and truncated Gaussian kernel function. In Figure 7, we present the 890 estimated densities.
Figure 7: A time series of estimated probability density functions via a kernel density estimator with truncated Gaussian kernel and Silverman’s ROT for the bandwidth selection. The estimated densities reflect daily COVID19 hospitalisation counts across the 96 departments.
From the estimated $(\widehat{f}_{1},\widehat{f}_{2},\dots,\widehat{f}_{890})$, we split the samples into an initial training sample and a testing sample. The initial training sample consists of the first 594 estimated densities, while the testing sample consists of the remaining 296 estimated densities. Using the first 594 estimated densities from March 18, 2020, to November 1, 2021, we produce one-step-ahead forecasts and evaluate its KLD with its holdout testing sample on November 2, 2021. Through an expanding window approach, we increase the training sample by one and use the first 595 estimated densities to produce one-step-ahead forecasts and evaluate its KLD on November 3, 2021. We proceed in the same manner to obtain 296 KLDs corresponding to the testing period.
In Figure 8, we present the 296 one-step-ahead density forecasts of the four methods considered. Since the regression mean function (i.e., signal) is unknown, the random walk is not considered a benchmark method. Compared with the holdout densities, we found the nonparametric density-on-density regression with the Bayes NW estimator performs the best. Similar to the CoDa method, it also uses the log-ratio transformation to remove constraints. A univariate functional time series forecasting method or a nonparametric function-on-function is then applied to obtain the forecasts. The forecasts are then back-transformed via the inverse log-ratio transformation. The nonparametric density-on-density regression with the Bayes NW estimator generally produces smoother density forecasts than the ones obtained from the CoDa method. The intuition is that the former one is a weighted average of the past densities and tends to be smooth.
Figure 8: For the 296 holdout densities, we present the one-step-ahead density forecasts of the four methods based on an expanding window.
In Table 2, we present summary statistics of the KLDs obtained from the four methods. Based on the median and mean values, the nonparametric density-on-density regression with the Bayes NW estimator performs the best among the four methods.
Table 2: From the 296 KLDs, we present summary statistics for the four methods.
          Statistic           Bayes NW           CoDa           LQDT           HZ
          Min.           0.0041           0.0051           0.0171           0.0176
          1st Qu.           0.0188           0.0197           1.0127           0.0727
          Median           0.0293           0.0324           4.1861           0.1747
          Mean           0.0420           0.0476           8.4794           0.2014
          3rd Qu.           0.0524           0.0554           12.8809           0.2910
          Max.           0.5241           0.4243           33.2268           0.9041
3.4 Analysis of age-specific life-table death counts in the United States
While it requires a way of estimating densities using a kernel estimator in Section 3.3, sometimes, we observe a time series of densities, such as period life table in demography and actuarial studies. A period life table is a table which shows, for each age, the probability that a person of that age will die before the next birthday. In the first age group, the initial number of alive is 100,000, while the remaining number of alive is 0 in the last age group. The sum of life-table death count is 100,000 every year, and the values are nonnegative, so these observations can be viewed as densities normalised to 100,000 rather than 1. By modelling the life-table death counts, we could understand a redistribution of survival probabilities, where deaths at younger ages gradually shift towards older ages. In Figure 1, we present rainbow plots of the female and male age distributions of life-table death counts in the USA from 1933 to 2023 in a single-year group.
Regarding the life-table death counts, we split the entire data set into a training sample from 1933 to 2000 and a testing sample from 2001 to 2023. Based on the initial training sample from 1933 to 2000, we compute the one-step-ahead density forecast in 2001 and the forecast error via the KLD. From the data in the training sample from 1933 to 2001, we again compute the one-step-ahead density forecast in 2002 and compute the forecast error. Then, we increase the training sample until the end of the testing samples. In Figure 9, we display the density forecasts computed by the nonparametric density-on-density regression with the Bayes NW estimator, CoDa method, LQDT method of Petersen & Müller (2016), and Horta & Ziegelmann’s (2018) method. Compared with the actual holdout samples, the Bayes NW estimator can adequately capture the local patterns, especially for the male life-table death counts between ages 18 and 40.
Figure 9: Holdout and forecast female and male life-table death counts from 2001 to 2023. The forecast life-table death counts were obtained via the nonparametric density-on-density regression with Bayes NW estimator, CoDa method, LQDT method, and HZ method, where the optimal bandwidth parameter was selected by the GCV.
With the optimal bandwidth parameter selected by the GCV, we iteratively compute the one-step-ahead density forecasts for years from 2001 to 2023, and then compute the overall KLDs in Figure 10.
Figure 10: Forecast accuracy (MSPE ratios) between the nonparametric density-on-density regression using the Bayes NW, CoDa method, LQD method, and HZ method for forecasting one-step-ahead age-specific life-table death counts from 2001 to 2023 in the USA.
We evaluate and compare the KLD among the Bayes NW, CoDa, LQDT and HZ methods. The KLD of the nonparametric density-on-density regression with the Bayes NW estimator is the smallest among the four methods considered. The inferior performance of the LQDT method is primarily due to the left boundary problem, where there is a comparably higher number of deaths at the infant age.
4 Conclusion
We propose a nonparametric density-on-density regression with the Bayes NW estimator to model and forecast density-valued objects. The advantage of the nonparametric density-on-density regression is that the functional time series forecasts within the CoDa do not require a dimension reduction. In turn, there is no loss of information in the proposed method. Simulation and empirical data analyses demonstrate that the nonparametric density-on-density regression achieves good forecast accuracy.
Via a series of simulation studies, we show that the proposed method is competitive with several existing methods reported in Kokoszka et al. (2019). Since the true densities are often unobserved in empirical applications, we consider a kernel density estimator to estimate densities. The estimation accuracy of such an estimator depends on the kernel function type and bandwidth parameter selection. For estimating densities, we consider the truncated Gaussian kernel function, and the optimal bandwidth is selected by either Silverman’s ROT. For estimating the nonparametric estimator of the density-on-density regression, the optimal bandwidth parameter is selected by the GCV. With the selected bandwidth, we evaluate and compare the one-step-ahead forecast error between forecast densities constructed based on observed and estimated densities based on a kernel density estimator of the holdout observations using the expanding window approach. For reproducibility, the [Uncaptioned image] code for the forecasting methods and their use in simulation and empirical data analyses are described in https://github.com/hanshang/Bayes_NW.
There are at least three ways the current paper can be further extended: First, a future extension is to develop a bandwidth procedure that jointly selects the optimal bandwidth parameters of the kernel density estimator and Bayes NW estimator. Second, comparing the predictive performance of different methods might not be as fruitful as one might think. It is hard to know when a predictive model works because everything is drowned in the incompressible randomness of new observations. The difference between competing methods can often be seen only in the digits of prediction error, even if they encode substantially different pictures of reality. One may achieve improved forecast accuracy by linearly combining the density forecasts from the nonparametric density-on-density regression and CoDa method. Finally, we proposed a Bayes NW estimator because of its simplicity. Its extension to Bayes local linear estimator can produce improved estimation accuracy.
Acknowledgments
The second author acknowledges financial support from an Australian Research Council Future Fellowship (grant no. FT240100338).
[…]
Appendix A Preliminaries
Conditional expectation in $B^{2}(I)$ and model. Let $g$ and $h$ two $\mathcal{B}^{2}(I)$-valued random PDFs; for any $f$ in $\mathcal{B}^{2}(I)$, the conditional expectation $\text{E}(g|h)$ is such that $\langle\text{E}(g|h),f\rangle_{\mathcal{B}}=\text{E}\left(\langle g,f\rangle_{\mathcal{B}}|h\right)$ (see, e.g., Bosq, 2000). Then, $\langle\text{E}(f_{t+1}|f_{t}),g\rangle_{\mathcal{B}}=\text{E}\left(\langle f_{t+1},\,g\rangle_{\mathcal{B}}|f_{t}\right)$ and (1) entails $\langle\text{E}(f_{t+1}|f_{t}),g\rangle_{\mathcal{B}}=\text{E}\left(\langle m(f_{t})\oplus\epsilon,\,g\rangle_{\mathcal{B}}|f_{t}\right)$. With the property of the inner product, $\langle\cdot,\,\cdot\rangle_{\mathcal{B}}$, $\langle\text{E}(f_{t+1}|f_{t}),g\rangle_{\mathcal{B}}=$ $\text{E}\left(\langle m(f_{t}),\,g\rangle_{\mathcal{B}}|f_{t}\right)+\text{E}\left(\langle\epsilon_{t},\,g\rangle_{\mathcal{B}}|f_{t}\right)=\langle m(f_{t}),\,g\rangle_{\mathcal{B}}+\langle\text{E}(\epsilon_{t}|f_{t}),\,g\rangle_{\mathcal{B}}$. According to Model (1), $\text{E}(\epsilon_{t}|f_{t})={\bf 0}_{\mathcal{B}}$ so that, for any $g$ in $\mathcal{B}^{2}(I)$, $\langle\text{E}(f_{t+1}|f_{t}),g\rangle_{\mathcal{B}}=\langle m(f_{t}),g\rangle_{\mathcal{B}}$, which is equivalent to $m(f)=\text{E}\left(f_{t+1}|f_{t}=f\right)$.
Comparing $\text{E}\left(\|\widehat{f}_{t}\ominus f_{t}\|_{\mathcal{B}}^{2}\right)$ and $\text{MISE}(\widehat{f}_{t})$. The next result gives a general situation where $\|f\ominus g\|_{\mathcal{B}}^{2}$ is smaller than $\|f-g\|$ up to a constant.
Proposition 1
Set $\theta>0$; for any $f,g\in\mathcal{B}^{2}(I)$ with $f,g\geq\theta$, $\|f\ominus g\|_{\mathcal{B}}^{2}=O(\|f-g\|^{2})$.
Proof. With the isometric feature of the clr transformation, $\|f\ominus g\|_{\mathcal{B}}^{2}=\int_{I}\{\text{clr}(f)-\text{clr}(g)\}^{2}$, and the definition of clr results in $\|f\ominus g\|_{\mathcal{B}}^{2}=\int_{I}(\log f-\log g)^{2}+\frac{3}{b-a}\left\{\int_{I}(\log f-\log g)\right\}^{2}$. According to the Lipschitz property of the $\log$ function on $[\theta,\,+\infty[$, $|\log f(u)-\log g(u)|\leq\theta^{-1}|f(u)-g(u)|$ as soon as $f(u)\geq\theta$ and $g(u)\geq\theta$ for any $u$; $\|f\ominus g\|_{\mathcal{B}}^{2}\leq M_{\theta}\,\|f-g\|^{2}$ where $M_{\theta}=\theta^{-2}\{1+3/(b-a)\}$. As a by-product, if $f$ is a density in $\mathcal{B}^{2}(I)$ such that $f\geq\theta>0$, then $\text{E}\left(\|\widehat{f}_{t}\ominus f_{t}\|_{\mathcal{B}}^{2}\right)=O\left\{\text{MISE}(\widehat{f}_{t})\right\}$.
Appendix B Proof of Theorem 1
Set $\widehat{K}_{t}=K_{r}\left(h^{-1}\,\|\widehat{f}_{t}\ominus f\|_{\mathcal{B}}\right)$, $\widehat{B}=(N\,\text{E}\widehat{K}_{1})^{-1}\sum^{N}_{t=1}\,\widehat{K}_{t}$, $\widehat{A}=(N\,\text{E}\widehat{K}_{1})^{-1}\odot\bigoplus_{t=1}^{N}\left(\widehat{K}_{t}\odot\widehat{f}_{t+1}\right)$; the proof is based on the decomposition
$\|\widehat{m}_{N}(f)\ominus m(f)\|_{\mathcal{B}}\leq\widehat{B}^{-1}\left\{Q_{1}+Q_{2}\right\}+\widehat{B}^{-1}\|m(f)\|_{\mathcal{B}}\,Q_{3},$ (4)
where $Q_{1}=\|\widehat{A}\ominus\text{E}\widehat{A}\|_{\mathcal{B}}$, $Q_{2}=\|\text{E}\widehat{A}\ominus m(f)\|_{\mathcal{B}}$, and $Q_{3}=|1-\widehat{B}|$. Before going on, let us point out useful features of the Bayes space $\mathcal{B}^{2}(I)$ properties that are systematically used in the proofs. First, the Bayes space of bounded PDFs is a separable Hilbert space; it exists an orthonormal basis $\{e_{j}\}_{j\geq 1}$ such that, for any $g$ in $\mathcal{B}^{2}(I)$, $g=\sum_{j\geq 1}\langle g,e_{j}\rangle_{\mathcal{B}}^{2}$. Second, the standard properties of norms naturally apply in $\mathcal{B}^{2}(I)$; for instance, for any $u_{1}$, $u_{2}$ in $\mathbb{R}$ and any $g_{1}$, $g_{2}$ in $\mathcal{B}^{2}(I)$, $\|u_{1}\odot g_{1}\oplus u_{2}\odot g_{2}\|_{\mathcal{B}}\leq|u_{1}|\|g_{1}\|_{\mathcal{B}}+|u_{2}|\|g_{2}\|_{\mathcal{B}}$. Third, with the assumptions on the kernel function (H(H11)), it exists $C>0$ and $C^{\prime}>0$ such that $C\,\pi_{f}(h_{r})\leq\text{E}\,K_{1}\leq C^{\prime}\pi_{f}(h_{r})$ (resp. $C\,\widehat{\pi}_{f}(h_{r})\leq\text{E}\widehat{K}_{1}\leq C^{\prime}\widehat{\pi}_{f}(h_{r})$), where $\pi_{f}(h_{r})=P\left(\|f_{1}\ominus f\|_{\mathcal{B}}\leq h_{r}\right)$ (resp. $\widehat{\pi}_{f}(h_{r})=P(\|\widehat{f}_{1}\ominus f\|_{\mathcal{B}}\leq h_{r})$). From now on, $C$ and $C^{\prime}$ denote two generic strictly positive constants.
Focus on $Q_{1}$. The separability property of the Bayes space $\mathcal{B}^{2}(I)$ entails the existence of an orthonormal basis $\{e_{j}\}_{j\geq 1}$ such that $\text{E}Q_{1}^{2}=\sum_{j\geq 1}\text{E}\left(\langle\widehat{A}\ominus\text{E}\widehat{A},e_{j}\rangle_{\mathcal{B}}^{2}\right)$. Let $u_{N}$ be a sequence tending to infinity with $N$: $\text{E}Q_{1}^{2}=Q_{1,1}+Q_{1,2}+Q_{1,3}$ with $Q_{1,1}=(N\,\text{E}\widehat{K}_{1})^{-2}\sum_{t=1}^{N}\sum_{j\geq 1}\text{Var}\left(\widehat{K}_{t}\langle\widehat{f}_{t+1},e_{j}\rangle_{\mathcal{B}}\right)$, $Q_{1,2}=(N\,\text{E}\widehat{K}_{1})^{-2}\underset{|s-t|<u_{N}}{\sum_{s=1}^{N}\sum_{t=1}^{N}}\sum_{j\geq 1}\text{cov}_{s,t,j}$, and $Q_{1,3}=(N\,\text{E}\widehat{K}_{1})^{-2}\underset{|s-t|>u_{N}}{\sum_{s=1}^{N}\sum_{t=1}^{N}}\sum_{j\geq 1}\text{cov}_{s,t,j}$, where $\text{cov}_{s,t,j}=\text{Cov}\left(\widehat{K}_{s}\langle\widehat{f}_{s+1},\,e_{j}\rangle_{\mathcal{B}},\,\widehat{K}_{t}\langle\widehat{f}_{t+1},e_{j}\rangle_{\mathcal{B}}\right)$. $Q_{1,1}\leq(N\,\text{E}\widehat{K}_{1})^{-2}\sum_{t=1}^{N}\text{E}\left(\widehat{K}_{t}^{2}\,\|\widehat{f}_{t+1}\|_{\mathcal{B}}^{2}\right)$ and $\|\widehat{f}_{t+1}\|_{\mathcal{B}}\leq\|\widehat{f}_{t+1}\ominus f_{t+1}\|_{\mathcal{B}}+\|f_{t+1}\ominus f\|_{\mathcal{B}}+\|f\|_{\mathcal{B}}$. (H(H5)) and (H(H11)) result in $\|\widehat{f}_{t+1}\|_{\mathcal{B}}=O_{P}(1)$; $Q_{1,1}=O\left(\left\{N\,\widehat{\pi}_{f}(h_{r})\right\}^{-1}\right)$. With the Cauchy-Schwartz inequality and the dependence model, $\sum_{j\geq 1}|\text{cov}_{s,t,j}|\leq 2\,\text{E}\left(\widehat{K}_{1}^{2}\|\widehat{f}_{2}\|_{\mathcal{B}}^{2}\right)$; again, with (H(H5)) and (H(H11)), $Q_{1,2}=O\left(u_{N}\left\{N\,\widehat{\pi}_{f}(h_{r})\right\}^{-1}\right)$. According to the definition of $\rho$-mixing sequence, $\underset{|s-t|>u_{N}}{\sum_{s=1}^{N}\sum_{t=1}^{N}}\sum_{j\geq 1}|\text{cov}_{s,t,j}|\leq\rho(u_{N})\,\text{E}\left(\widehat{K}_{1}^{2}\|\widehat{f}_{2}\|_{\mathcal{B}}^{2}\right)$; then (H(H3)) results in $Q_{1,3}=O\left(u_{N}^{-a}\,\widehat{\pi}_{f}(h_{r})^{-1}\right)$. If we set $u_{N}=N^{1/(a+1)}$, the quantity balancing $Q_{1,2}$ and $Q_{1,3}$, then $Q_{1}=O_{P}\left(\left\{N^{a/(a+1)}\,\widehat{\pi}_{f}(h_{r})\right\}^{-1/2}\right)$.
Focus on $Q_{2}$. Taking into account the dependence model, $Q_{2}=Q_{2,1}+Q_{2,2}+Q_{2,3}$ with $Q_{2,1}=(\text{E}\widehat{K}_{1})^{-1}\text{E}\left(\widehat{K}_{1}\|\widehat{f}_{2}\ominus f_{2}\|_{\mathcal{B}}\right)$, $Q_{2,2}=(\text{E}\widehat{K}_{1})^{-1}\text{E}\left(\widehat{K}_{1}\|m(f_{1})\ominus m(f)\|_{\mathcal{B}}\right)$, $Q_{2,3}=(\text{E}\widehat{K}_{1})^{-1}\left\|\text{E}\left(\widehat{K}_{1}\odot\epsilon_{1}\right)\right\|_{\mathcal{B}}$. Cauchy-Schwartz inequality with (H(H5)) and (H(H11)) entails $\text{E}\left(\widehat{K}_{1}\|\widehat{f}_{2}\ominus f_{2}\|_{\mathcal{B}}\right)=O\left(\widehat{\pi}_{f}(h_{r})^{1/2}\delta_{N}\right)$; use again (H(H11)) to get $Q_{2,1}=O\left(\widehat{\pi}_{f}(h_{r})^{-1/2}\delta_{N}\right)$. With the Lipschitz property of the regression operator $m$ (see (H(H1))), $\|m(f_{1})\ominus m(f)\|_{\mathcal{B}}\leq C\|f_{1}\ominus f\|_{\mathcal{B}}$. Because we only retain $f_{t}$’s such that $K_{t}>0$, $\|m(f_{1})\ominus m(f)\|_{\mathcal{B}}\leq C\,h$ and $Q_{2,2}=O\left(h_{r}\right)$. $\text{E}\left(K_{1}\odot\epsilon_{1}\right)=\text{E}\left\{\text{E}\left(K_{1}\odot\epsilon_{1}|f_{1}\right)\right\}$: $\forall g\in\mathcal{B}^{2}(I)$, $\langle\text{E}\left(K_{1}\odot\epsilon_{1}|f_{1}\right),\,g\rangle_{\mathcal{B}}=K_{1}\,\langle\text{E}\left(\epsilon_{1}|f_{1}\right),\,g\rangle_{\mathcal{B}}=K_{1}\,\langle{\bf 0}_{\mathcal{B}},\,g\rangle_{\mathcal{B}}=0$. Conclusion: for any $g$ in $\mathcal{B}^{2}(I)$, $\langle\text{E}\left(K_{1}\odot\epsilon_{1}|f_{1}\right),\,g\rangle_{\mathcal{B}}=0$, which results in $Q_{2,3}=0$. All these intermediate results in $Q_{2}=O(h_{r})+O\left(\widehat{\pi}_{f}(h_{r})^{-1/2}\delta_{N}\right)$.
Remaining terms and summary. $Q_{3}$ is a particular case of $Q_{1}$ when $\widehat{f}_{t+1}\equiv 1$ and $\widehat{B}=1+(\widehat{B}-1)$ entails $\widehat{B}\leq 1+Q_{3}$ so that $\widehat{B}=1+O_{P}\left(\left\{N^{a/(a+1)}\,\widehat{\pi}_{f}(h_{r})\right\}^{-1/2}\right)$. Then, the decomposition (4) results in
$\|\widehat{m}_{N}(f)\ominus m(f)\|_{\mathcal{B}}=O(h_{r})+O\left(\widehat{\pi}_{f}(h_{r})^{-1/2}\delta_{N}\right)+O_{P}\left(\left\{N^{a/(a+1)}\,\widehat{\pi}_{f}(h_{r})\right\}^{-1/2}\right).$ (5)
[…]
To end the proof of Theorem 1, just replace $\widehat{\pi}_{f}(h_{r})$ with $\pi_{f}(h_{r})$ into (5).